Set up an objective of your portfolio (high risk, low risk, target audience , benchmark ) objective should not exceed 250 words. 

 

4. Prepare returns for assets, compute beta, alpha of individual stock and protfolio as whole

 

5. Use solver function in excel to maximise returns and minimise risks.

 

6. Draw efficient market frontier of your data. 

 

7. Assuming you have held this portfolio for an year now. Comment on last years performance of your portfolio against benchmark (500 words maximum), write economic analysis (In this case USA if yu have selected companies from US Stock exchange) 250 words maximum along side industry analysis (250 words)

 

8. Students with highest risk and return ratio will be awarded an extra 2 marks as cherry on top for his stock selection.


1
Expert's answer
2021-06-02T16:54:13-0400

3) The risk portfolio objective is to help an individual and the financial advisor to choose the right mix of investment. The high risk investment have 70% and above equities while low risk has below 40% equities.

4) return on assets

A portfolio has three investment A,B,C with weights of 35%,25% and 40% respectively.Their expected returns are 6% ,7%, 10%respectively.

The expected return of portfolio will be

"(35×6)+(25×7)+(40×10)=7.85\%"

Computation of beta in a portfolio

"Beta=\frac{covariance}{variance}"

Given the correlation coefficient between the company and market returns as 0.2 and standard deviation of the return of the company and market as 6 % and 4% respectively, the beta will be

"=\frac{0.2×6}{4}=0.3"

Computation of alpha in a portfolio

Alpha=portfolio return-risk free return-(beta ×market return)

"=7.85-4-(0.3×0.2)=3.79\%"

5)


6)


7) The investment portfolio is satisfying since the objective has been achieved with proper assets allocation, tax effectiveness low expense ratio.